# Strategic Informed Trading and the Value of Private Information

### v1

@@ Appendix E Proofs from Section 3
$\begin{split}&\gamma_{U}\textrm{CE}^{U}_{0,\iota}(H_{\iota})=\psi_{U,0}^{\prime}\mu_{X}+\psi_{U,0}^{\prime}P_{X}^{-1/2}\bigg{(}\frac{\widehat{y}}{(1+p_{I})(1+2\widehat{y})}\left(p_{I}\mathcal{E}_{X}+\sqrt{p_{I}}\mathcal{E}_{I}+\frac{\widehat{y}+1}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}\right)\\ &\quad+\frac{\widehat{y}\lambda(\kappa p_{I}+(1+\widehat{y})^{2})}{(1-\lambda)(1+2\widehat{y})(\widehat{y}+1)^{2}}P_{X}^{-1/2}\psi_{I,0}-\frac{\kappa p_{I}+(1+\widehat{y})^{2}}{(1-\lambda)(1+2\widehat{y})(1+\widehat{y})}P_{X}^{-1/2}\widehat{\Pi}\bigg{)}\\ &\quad+\frac{1}{2}\frac{\lambda^{2}(\kappa p_{I}+(1+\widehat{y})^{2})}{(1-\lambda)^{2}(1+p_{I})(1+2\widehat{y})^{2}\left(\kappa p_{I}+\frac{1}{1+p_{I}}(1+\widehat{y})^{2}\right)}\bigg{\|}p_{I}\mathcal{E}_{X}+\sqrt{p_{I}}\mathcal{E}_{I}+\frac{\widehat{y}+1}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}\\ &\qquad\qquad+\frac{\left(\kappa p_{I}+\frac{1}{1+p_{I}}(1+\widehat{y})^{2}\right)\widehat{y}(1+p_{I})}{(1+\widehat{y})^{2}}P_{X}^{-1/2}\psi_{I,0}-\frac{\left(\kappa p_{I}+\frac{1}{1+p_{I}}(1+\widehat{y})^{2}\right)(1+p_{I})}{(1+\widehat{y})\lambda}P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
Using $\widehat{y}\to\lambda/(1-\lambda)$ in the limit, we obtain the almost sure limit
@@
$\begin{split}&\lim_{p_{I}\to 0}\gamma_{U}\textrm{CE}^{U}_{0,\iota}(H_{\iota})=\psi_{U,0}^{\prime}\mu_{X}+\frac{1}{1-\lambda^{2}}\psi_{U,0}^{\prime}P_{X}^{-1/2}\bigg{(}\frac{\lambda}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}+\lambda^{2}P_{X}^{-1/2}\psi_{I,0}\\ &\qquad-P_{X}^{-1/2}\widehat{\Pi}\bigg{)}+\frac{1}{2(1-\lambda^{2})^{2}}\bigg{\|}\frac{\lambda}{\alpha_{I}\sqrt{p_{N}}}\mathcal{E}_{N}+\lambda^{2}P_{X}^{-1/2}\psi_{I,0}-P_{X}^{-1/2}\widehat{\Pi}\bigg{\|}^{2}.\end{split}$
Substituting back in $Z_{N}$ and plugging in for $\widehat{\Pi}$, the right side is
$\begin{split}&\psi_{U,0}^{\prime}\mu_{X}\pm\frac{1}{2}\psi_{U,0}^{\prime}P_{X}^{-1}\psi_{U,0}+\frac{1}{\lambda_{U}(2-\lambda_{U})}\psi_{U,0}^{\prime}P_{X}^{-1}\bigg{(}\frac{\lambda_{I}}{\alpha_{I}}Z_{N}-\lambda_{U}\lambda_{I}\psi_{I,0}\\ &\qquad-\lambda_{U}\psi_{U,0}\bigg{)}+\frac{1}{2\lambda_{U}^{2}(2-\lambda_{U})^{2}}\bigg{\|}P_{X}^{-1/2}\left(\frac{\lambda_{I}}{\alpha_{I}}Z_{N}-\lambda_{U}\lambda_{I}\psi_{I,0}-\lambda_{U}\psi_{U,0}\right)\bigg{\|}^{2},\end{split}$
from which the result follows. ∎

### v2

@@ B.5. Proofs from Section 3
$\begin{split}\textrm{CE}^{U}_{0}(H_{\iota})&=\textrm{CE}^{o}_{U,0}(H_{\iota})+\frac{\alpha_{U}\lambda_{I}^{2}R(1+\beta)}{2(\lambda_{I}+\beta+R\lambda_{U})^{2}(R+\beta)}\times\left|(1-R)P^{1/2}_{X}(H-p_{{\mathbb{Q}}_{0}})\right|^{2},\\ \textrm{CE}^{U}_{0,\iota}(H_{\iota})&=\textrm{CE}^{o}_{U,0}(H_{\iota})\\ &\qquad+\frac{\alpha_{U}\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})}{2\lambda_{U}^{2}R(\beta+R(1+\widehat{y})^{2})(1+2\widehat{y})^{2}}\times\left|(1-R)P^{1/2}_{X}(H_{\iota}-p_{{\mathbb{Q}}_{0}})\right|^{2},\end{split}$
where
@@
$\begin{split}(1-R)H_{0}&=(1-R)X_{0}+\sqrt{R(1-R)}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{R}{\alpha_{I}}P_{X}^{-1}Z_{N},\\ (1-R)H_{0,\iota}&=(1-R)X_{0}+\sqrt{R(1-R)}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{(1+\widehat{y})R}{\alpha_{I}}P_{X}^{-1}Z_{N}.\end{split}$
The results for $U$ readily follow.